# Vogelpothsweg 87, D-44227 Dortmund, Germany

### v2

@@ 2 Element-based Rusanov method
We assume that periodic boundary conditions are imposed on $\partial\Omega$ and that there exists a convex invariant set $\mathcal{G}$ such that $u\in\mathcal{G}$ a.e. in $\Omega\times(0,T]$. In the context of gas dynamics, all states belonging to $\mathcal{G}$ have positive density and internal energy.
A continuous piecewise-linear ($\mathbb{P}_{1}$) finite element approximation $u_{h}=\sum_{j=1}^{N_{h}}u_{j}\varphi_{j}$ to $u$ is called invariant domain preserving (IDP) or positivity preserving if $u_{j}\in\mathcal{G}$ for $j=1,\dots,N_{h}$. The Lagrange basis functions $\varphi_{i}$ are associated with vertices $\mathbf{x}_{i}$ of a conforming mesh $\mathcal{T}_{h}$. The elements (cells) of that mesh are denoted by $E_{1},\ldots,E_{h}$. The global indices of nodal points $\mathbf{x}_{i}$ belonging to $K_{e}$ are stored in the set $\mathcal{N}^{e}$. The indices of elements $K_{e}$ containing $\mathbf{x}_{i}$ are stored in the set $\mathcal{E}_{i}$.
@@
The element-based low-order method (3) is a slightly modified version of the one derived in (book, , Sec. 4.6.1) using a trapezoidal rule approximation. If (3) is discretized in time using an explicit strong stability preserving (SSP) Runge–Kutta method, then each low-order forward Euler stage can be written as
@@ 3 Residual distribution
where $\mathbf{n}$ denotes the unit outward normal to $\partial K_{e}$. The general form (5) with
$r_{i}^{e}=r_{i}^{e,L}+f_{i}^{e,*}$
reduces to (2) for $f_{i}^{e,*}=f_{i}^{e}$ and to (3) for $f_{i}^{e,*}=0$. The fluctuation $r^{e}=\sum_{i\in\mathcal{N}^{e}}r_{i}^{e}$ defined in formula (6) is preserved provided that $\sum_{i\in\mathcal{N}^{e}}f_{i}^{e,*}=0$.
@@ 4.1 Limiting for scalars
    $f_{i}^{e,*}=\begin{cases}-(\tilde{f}^{e}_{-}/\tilde{f}^{e}_{+})\tilde{f}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}>0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}>0,\\ -(\tilde{f}^{e}_{+}/\tilde{f}^{e}_{-})\tilde{f}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}<0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}<0,\\ \tilde{f}_{i}^{e}&\mbox{otherwise}.\end{cases}$
@@
Note that $\tilde{f}_{i}^{e}=\tilde{\alpha}_{i}^{e}f_{i}^{e}$, where $\tilde{\alpha}_{i}^{e}$ corresponds to $\alpha_{i}^{e}$ defined by (9). Hence, the result of C&S limiting can be written as $f_{i}^{e,*}=\alpha_{i}^{e}f_{i}^{e}$, where
$\alpha_{i}^{e}=\begin{cases}-(\tilde{f}^{e}_{-}/\tilde{f}^{e}_{+})\tilde{\alpha}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}>0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}>0,\\ -(\tilde{f}^{e}_{+}/\tilde{f}^{e}_{-})\tilde{\alpha}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}<0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}<0,\\ \tilde{\alpha}_{i}^{e}&\text{otherwise}.\end{cases}$
Further examples of scalar element-based limiters can be found in (book, , Chap. 4).
@@ 4.2 Limiting for systems
Let us now consider a system of conservation laws for $u=(\varrho,\varrho\phi_{2},\ldots,\varrho\phi_{m})^{\top}$, where $\varrho=\rho\phi_{1}$ with $\phi_{1}=1$ is a density-like variable. The invariant set to be preserved by an IDP limiter for the element contributions $f_{i}^{e}=(f_{i,\varrho}^{e},f_{i,\varrho\phi_{1}}^{e},\ldots,f_{i,\varrho\phi_{m}}^{e})^{\top}$ is usually of the form $\mathcal{G}=\{u\in\mathbb{R}^{m}\,:\,\Phi_{1}(u)\geq 0,\ldots,\Phi_{M}(u)\geq 0\}$, where $\Phi_{1},\ldots,\Phi_{M}$ are quasi-concave scalar functions of the conserved variables.

### v3

@@ 2 Element-based Rusanov method
We assume that periodic boundary conditions are imposed on $\partial\Omega$ and that there exists a convex invariant set $\mathcal{G}$ such that $u\in\mathcal{G}$ a.e. in $\Omega\times(0,T]$.
A continuous piecewise-linear ($\mathbb{P}_{1}$) finite element approximation $u_{h}=\sum_{j=1}^{N_{h}}u_{j}\varphi_{j}$ to $u$ is called invariant domain preserving (IDP) or positivity preserving if $u_{j}\in\mathcal{G}$ for $j=1,\dots,N_{h}$. The Lagrange basis functions $\varphi_{i}$ are associated with vertices $\mathbf{x}_{i}$ of a conforming mesh $\mathcal{T}_{h}$. The elements (cells) of that mesh are denoted by $K_{1},\ldots,K_{E_{h}}$. The global indices of nodal points $\mathbf{x}_{i}$ belonging to $K_{e}$ are stored in the set $\mathcal{N}^{e}$. The indices of elements $K_{e}$ containing $\mathbf{x}_{i}$ are stored in the set $\mathcal{E}_{i}$.
@@
If (3) is discretized in time using an explicit strong stability preserving (SSP) Runge–Kutta method, then each low-order forward Euler stage can be written as
@@
The element-based low-order method (3) is new but represents a slightly modified version of the one derived in (book, , Sec. 4.6.1) using a trapezoidal rule approximation for the flux divergence term. The use of the midpoint rule reduces the effort for calculating the Rusanov viscosity $d^{e}$ and makes it possible to prove the IDP property without introducing edge-based bar states (cf. (book, , Thm. 4.13)).
@@ 3 Residual distribution
where $\mathbf{n}$ is the unit outward normal. The general form (5) with $r_{i}^{e}=r_{i}^{e,L}+f_{i}^{e,*}$ reduces to (2) for $f_{i}^{e,*}=f_{i}^{e}$ and to (3) for $f_{i}^{e,*}=0$. The fluctuation $r^{e}=\sum_{i\in\mathcal{N}^{e}}r_{i}^{e}$ defined in formula (6) is preserved provided that $\sum_{i\in\mathcal{N}^{e}}f_{i}^{e,*}=0$.
@@ 4.1 Limiting for scalars
    $f_{i}^{e,*}=\begin{cases}-(\tilde{f}^{e}_{-}/\tilde{f}^{e}_{+})\tilde{f}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}>0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}>0,\\ -(\tilde{f}^{e}_{+}/\tilde{f}^{e}_{-})\tilde{f}_{i}^{e}&\mbox{if}\ \tilde{f}_{i}^{e}<0,\ \tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}<0,\\ \tilde{f}_{i}^{e}&\mbox{otherwise}.\end{cases}$ (10)
@@
Note that $\tilde{f}_{i}^{e}=\tilde{\alpha}_{i}^{e}f_{i}^{e}$, where $\tilde{\alpha}_{i}^{e}$ corresponds to $\alpha_{i}^{e}$ defined by (9). Hence, the result of C&S limiting can be written as $f_{i}^{e,*}=\alpha_{i}^{e}f_{i}^{e}$, where $\alpha_{i}^{e}=-(\tilde{f}^{e}_{\mp}/\tilde{f}^{e}_{\pm})\tilde{\alpha}_{i}^{e}$ if $\tilde{f}^{e}_{+}+\tilde{f}^{e}_{-}\neq 0$. That is, the C&S algorithm applies an individually chosen correction factor $\alpha_{i}^{e}$ to each component $f_{i}^{e}$. As shown in lohmann2017 , this nodal limiting strategy is less diffusive than simple scaling using the common correction factor (9). Moreover, the limited element contributions (10) depend continuously on the data.
Further examples of scalar element-based limiting techniques can be found in (book, , Chap. 4). A classical representative of RD-based nodal limiters is the nonlinear positive streamwise invariant (PSI) method abgrall2006 ; RD1 ; struijs1994 . Such limiters produce distribution weights $\beta_{i,\pm}^{e}$ instead of correction factors $\alpha_{i}^{e}$. We remark that, while representations of $f_{i}^{e,*}$ in terms of $\alpha_{i}^{e}$ or $\beta_{i,\pm}^{e}$ may be useful for derivation and comparison purposes, direct computation of $f_{i}^{e,*}$ should be preferred in practice.
@@ 4.2 Limiting for systems
Let us now consider a system of conservation laws for $u=(\varrho,\varrho\phi_{2},\ldots,\varrho\phi_{m})^{\top}$, where $\varrho=\rho\phi_{1}$ with $\phi_{1}=1$ is a density-like variable. The invariant set to be preserved by an IDP limiter for the element contributions $f_{i}^{e}=(f_{i,\varrho}^{e},f_{i,\varrho\phi_{1}}^{e},\ldots,f_{i,\varrho\phi_{m}}^{e})^{\top}$ is usually of the form $\mathcal{G}=\{u\in\mathbb{R}^{m}\,:\,\Phi_{1}(u)\geq 0,\ldots,\Phi_{M}(u)\geq 0\}$, where $\Phi_{1},\ldots,\Phi_{M}$ are quasi-concave scalar functions of the conserved variables. The quasi-concavity implies that $\Phi(u)\geq 0$ for any convex combination of intermediate states $\bar{u}$ such that $\Phi(\bar{u})\geq 0$. This property is exploited in low-order IDP schemes and in limiting techniques based on convex decompositions guermond2016 ; convex ; book ; wu2025 .